A volume of 150 m3 of room air contains 1800 g at 22 ° C water vapor. What is the relative humidity of room air? Pressure 101.3 kPa.
1
Expert's answer
2019-03-13T15:16:18-0400

Solution:

Relative humidity is the ratio of vapor density to saturation vapor density:


"\phi=\frac{\rho} {\rho_0} 100\%"

Vapor density is equal to mass of vapor divided by the volume of the room:


"\rho=\frac{m} {V}"


Therefore:

"\phi=\frac{m} {V\rho_0} 100\%"

Saturation vapor density at "22^\circ\:C:"

"\rho_0=19.4\:g/m^3"

Now we can calculate relative humidity:

"\phi=\frac{1800} {150\times19.4} 100\%=62\%"

Answer:

"\phi=62\%"